# Reconstructing Sepsis Trajectories from Clinical Case Reports using LLMs: the Textual Time Series Corpus for Sepsis

### v1

@@ 1 Introduction
The Third International Consensus Definitions for Sepsis (Sepsis-3) are the benchmark definitions for sepsis, the dysfunctional immune response to an infection, which are used in trial eligibility criteria (Kyriazopoulou et al., 2021) and phenotyping studies (Seymour et al., 2019). In the critical care literature, sepsis is seen as a heterogeneous disease, a pathway arrived upon from a multitude of infectious origins, anatomical sites, and comorbid profiles. There is increasing recognition of the importance of time in understanding sepsis and sepsis progression, as seen in early-warning systems (Henry et al., 2022), variation in prediction utility (Kamran et al., 2024), and endotype characterization (Noroozizadeh et al., 2023).
Meanwhile, structured data streams, which are the data resources these tools use, are often incomplete records with insufficient information to render or confirm a diagnosis of sepsis (Moldwin et al., 2021). A more complete alternative is the discharge summary, or more generally, the case report, in which a clinician documents the clinical findings pertinent to the case and care management. The impracticality of using the case report is that it is only written in full after the encounter is over, i.e., with a less granular timestamp than the events contained within. In principle, having a more complete, fine-grained temporal record could enable better temporal analyses of patients with sepsis, leading to better understanding and treatment improvements.
At present, the (semi-)public critical care repositories at the research frontier contain high resolution structured signals and low resolution textual signals. While several concerted efforts have focused on extracting temporal information, they have largely focused on temporal relations between clinical concepts (Sun et al., 2013) rather than on event timing. Those that have focused on timing (Leeuwenberg and Moens, 2020; Frattallone-Llado et al., 2024) used excerpted summaries and relatively small sample sizes.
In order to train sepsis models that can consider sources of information references in free-text and that have fine time-granularity, we seek to generate larger samples of textual time series. To do so, we introduce a novel Sepsis-3 textual time series corpus from the Pubmed Open Access (PMOA) Subset. We adopt the LLM-as-annotator approach, evaluating its ability to extract clinical findings and its ability to accurately associate them with timestamps. We augment existing clinical annotations with our own on PMOA, and contrast LLM annotations with our own and that of Frattallone et al. (2024). To provide checks on their performance at scale, we consider a larger subsample for inter-LLM annotator agreement, and show the performance levels across several model sizes, which provides valuable context for cost-based analyses. Finally, these assessments provide evidence of the quality of the LLM annotations and our pipeline, which are then applied to generate the textual time series corpus for sepsis (T2S2).
@@ 1.1 Related work
Numerous critical care and sepsis datasets are available (semi-)publicly, e.g., Thoral et al. (2021); Pollard et al. (2018), including several with de-identified clinical discharge summaries: MIMIC-III (Johnson et al., 2016) and MIMIC-IV (Johnson et al., 2023). Sepsis phenotyping (including Sepsis-3) has been conducted using a variety of methods on structured data (Johnson et al., 2018), although structured data is often highly missing, leading to suboptimal phenotypes (Seinen et al., 2025). Additionally, the timing of sepsis phenotyping is crucial, in that delays in sepsis prediction can nullify predictiveness and lead to burdensome and unhelpful analytic models (Kamran et al., 2024).
Several works have used the i2b2 (a competition subset of MIMIC-II/III) to construct timelines of clinical concepts (Leeuwenberg and Moens, 2020; Frattallone-Llado et al., 2024). In the case of Leeuwenberg and Moens (2020), the event spans are already given, and focus on an excerpt of the full discharge summary, so we focus our direct comparison on that of Frattallone-Llado et al. (2024). In contrast to these works, we adopt an annotation process focused on temporal assignment to clinical finding rather than i2b2 clinical concept (Uzuner et al., 2011), which allows for greater specificity of the event in the absence of additional context. Additionally, as compared to Frattallone-Llado et al. (2024), we focus on using text alone for clinical time series construction, which is necessary for the Pubmed Open-Access Subset since no complementary data source is available.
Broadly in the medical domain, LLMs are showing promise in medical tasks such as summarization (Van Veen et al., 2024). In fact, some findings suggest that medical fine-tuning rarely leads to improved model over their corresponding foundation models (Jeong et al., 2024). In our analysis, we primarily use foundation LLMs and show that their scale and instruction-tuning are central to the performance of the model.
@@ 1.2 Contributions
1.3 Generalizable Insights
Access to the best temporal information available is critical in sepsis prediction tasks and has major implications in clinical decision making and early warning. The analysis pipeline we validate and corpus we introduce enable the use of textual data with temporal granularity, expanding the available features for use in a multitude of tasks. For example, the recent investigation in JAMA on vasopressin treatment policies using offline reinforcement learning uses hourly structured data in its learning, its reweighting analyses, and its validations (Kalimouttou et al., 2025). Access to textual features would augment the information that could affect its predictions, recommendations, and its adjustment for measured confounding, which may alter and refine its main findings. By adopting or extending our approach, the temporal resolution of the clinical event findings available only in text could be used and accounted for. As observed in Seinen et al. (2025), this may be a large fraction of concepts, with only 13% of textual events identified in structured form in their study.
@@ 2 Methods
Our task is the following: given a biomedical free-text corpus, construct a textual time series case report corpus of patients with Sepsis-3. We define the term textual time series as a list of clinical findings each with an associated timestamp (which may be absolute or relative to time of case presentation) pertaining to an individual. A clinical finding is a free-text specification of an entity pertaining to or with the potential to affect the person’s health.
@@
Figure 1: Flowchart of the PMOA T2S2 corpus construction pipeline (left). Confusion matrices (right) for Sepsis-3 phenotype in i2m4 (top) and sepsis-100 (bottom) datasets, with preferred annotator in column position.
Compared to clinical concept annotations (Uzuner et al., 2011; Sun et al., 2013), we choose to expand the text span determination to increase the specificity of the meaning attached to the finding. We note the following distinctions of our notion of clinical finding with that of i2b2 clinical concept, as per the i2b2 guideline specifications (Uzuner et al., 2011):
@@
The other design choice we make is to alter the time expression to disambiguate timing by capturing the start time of the perception of the clinical event. For example, if a patient has a “three-day history of fever”, the clinical event “fever” is assigned to -72 hours, since it is first perceived then. If the patient has a “history of smoking”, the start time is fully ambiguous, so the time of perception is the time that smoking is assessed, e.g., at the start of the encounter, and the event is “history of smoking”, rather than “smoking”. To restate, our clinical finding is defined as a contiguous text span from the text corresponding to a temporally localizable event pertaining to the patient, modified for conjunction splitting and temporal perception as above.
@@ 2.1 Data
We used two data sources: (1) the absolute timeline annotations from Frattallone-Llado et al. (2024), and (2) the Pubmed Open Access subset (PMC Open Access Subset, 2024) (see Table 1).
@@
The absolute timeline annotations (Frattallone-Llado et al., 2024) came from physician annotations of 20 discharge summaries following the i2b2 concept and annotation guidelines (note, the original competition annotation focused on discharge summary excerpts rather the entire note), with the addition of interval timestamping of the identified concepts. For our analysis, we used the lower bound of the interval (specified as a mean time) as it denotes the time when the finding may first be observed. We used only the probabilistic annotations, since the bounded annotation lower bound could precede the actual occurrence of the finding. Fifteen of the summaries come from i2b2 and five are from MIMIC IV (annotated in the same way)—hence we call this dataset “i2m4”.
Table 1: Dataset summary. Bold cells indicate contributed labeling
Corpus i2m4 Sepsis-10   Sepsis-100
N (unlabeled) 310; 331,794 2,139 2,139
N (labeled) 20 10 100
Source i2b2; MIMIC IV PMOA PMOA
Note type discharge summary case report case report
Labeler clinicians clinicians O1-preview
Phenotyper clinicians clinicians QwQ-32B-Preview
Methods O1-preview O1-preview
DeepSeek R1 IQ1 DeepSeek R1 IQ1 DeepSeek R1 IQ1
L3.3 70B Instruct L3.3 70B Instruct L3.3 70B Instruct
L3.1 70B Instruct L3.1 70B Instruct L3.1 70B Instruct
Mixtral 8x7B Mixtral 8x7B Mixtral 8x7B
 Instruct v0.1  Instruct v0.1  Instruct v0.1
L3.1 8B Instruct L3.1 8B Instruct L3.1 8B Instruct
Ablations No role No role No role
Base: No conjunction No conjunction No conjunction
L3.3 70B Instruct  instruction  instruction  instruction
No few-shot prompting No few-shot prompting No few-shot prompting
Sensitivity analyses With upper bound With upper bound With upper bound
With upper bound + With upper bound + With upper bound +
 i2b2 event typing  i2b2 event typing  i2b2 event typing
@@ 2.1.2 PMOA (sepsis-10, sepsis-100) Dataset
The second dataset comes from the Pubmed Open Access Subset, an open-access, free-text corpus of roughly 1.5 million biomedical publications (PMC Open Access Subset, 2024). Because the majority of these publications are not case reports, we built an extraction pipeline to recover critical care case reports, i.e., a clinical research informatics critical care cohort.
We used the PMOA Subset (version as of April 25, 2024). We removed the header and footer of the free-text which contained the Title, Abstract, and References sections, and retained the ==== Body section. To identify potential case reports, we required that the free-text match the case-insensitive regular expressions case (report|presenta) and year-? ?old. We contrasted the output of the query with PubMed metadata for case reports and found our regular expression approach had substantially higher specificity during inspection of a sample of 20 random reports.
To identify potential sepsis case reports, we filtered the potential case reports using the case-insensitive regular expressions (sepsi||||septic) and (critical||||intensive) care. Then, we used the an LLM query to extract sepsis cases, as well as the number of cases present, and the age and gender. We used GPT-4o and Llama-3.1-8B-Instruct for the phenotyping query screen. Reports were omitted if the number of cases identified was not equal to 1. As a screen, we chose to include the case report if either GPT-4o or Llama-3.1-8B-Instruct identified the report as having Sepsis-3.
@@ 2.2 Textual Time Series Annotations
18 years old |||| 0
male |||| 0
admitted to the hospital |||| 0
fever |||| -72
rash |||| -72
acne |||| -672
minocycline |||| -672
increased WBC count |||| 0
eosinophilia|||| 0
systemic involvement|||| 0
diffuse erythematous or maculopapular eruption|||| 0
pruritis |||| 0
DRESS syndrome |||| 0
fever persisted |||| 0
rash persisted |||| 0
discharged |||| 24
@@
For the i2m4 dataset, we also compared against a BERT-based baseline (BERT 5-CV), our adaption of the Frattallone et al. model (2024). In their work, they transformed the temporal prediction task into multiclass classification and used a multilayer perceptron (MLP) head with a softmax and cross-entropy loss for training. To predict relative time (hours from case presentation), we made the following adaptations. Similarly to the classifier module, we used a 3-layer MLP regressor head from the output of the BERT event span features and output the prediction means and standard deviations. The regression task (predicting the lower bound) was optimized using L1 loss. We added the BERT features SUBJ_START and SUBJ_END tokens surrounding the event span (Zhong and Chen, 2021) and removed the dropout layers inside the MLP heads, leaving only one at the final layer. Regarding the timestamps used for prediction, while many events occurred near the admission time, a subset extended months or years away, complicating the design of a regressor capable of both spanning wide magnitudes and maintaining precision for the majority of the means close to the admission time. To address this, we filtered the findings by their magnitudes, retaining them if the means were within two weeks of the admission time and the standard deviations were under two weeks. Filtering reduced the dataset size (in number of findings) from 3,472 to 3,350. Training was carried out for 50 epochs with batch size 32 and learning rate 3e-5. The model training used 5-fold cross-validation.
@@ 2.3 Assessments
We assessed the textual time series with the following methods: event match rate, concordance (c-index), median absolute time discrepancy, and log-time discrepancy. To determine event matches between two lists of strings, one for the label clinical findings and one for the predicted clinical findings, we used a recursive best match procedure, where a match is identified by the distance between two strings (see Appendix B for pseudocode). Specifically, the recursive best match was used to reduce erroneous surjective matches, i.e., multiple events from the predicted event list mapping to the same event in the reference list. For each reference event, if multiple predicted events matched, the lowest distance match was selected (with event ordering used to break distance ties), and both the predicted and referent events were removed from the lists for the recursive call to match events. Then all matched events from the recursive calls were returned. To compute distance, we tried (1) Levenshtein distance, (2) cosine distance between mean sentence embeddings of BERT (bert-base-uncased) and (3) cosine distance between mean sentence embeddings of PubMedBert (S-PubMedBert-MS-MARCO) (Deka et al., 2022). Upon manual review of event matches, we selected option 3 with a cosine similarity cutoff of 0.1.
@@ 2.3.1 Log-Time Cumulative Distribution Function
We define the log-time cumulative distribution function as follows. For each matched event $i$, compute the discrepancy $s_{i}=|t^{p}_{i}-t^{r}_{i}|$. then truncate at the maximum cutoff $S_{\text{max}}$ (e.g., 1 year = 8760 hours):
$s^{\prime}_{i}=\min(s_{i},S_{\text{max}})$
Apply the log transformation:
$x_{i}=\log(1+s^{\prime}_{i})$
Sort $x_{i}$ in ascending order:
$x_{(1)}\leq x_{(2)}\leq\cdots\leq x_{(k)}$
The CDF $F(x)$ is defined as the fraction of log-transformed discrepancies less than or equal to $x$:
$F(x)=\frac{1}{k}\sum_{i=1}^{k}\mathbf{1}_{\{x_{(i)}\leq x\}}$
where $\mathbf{1}$ is the indicator function.
Then, we get the following boundary values: $F(0)\geq 0$ (depending on the smallest discrepancy) and $F(\log(1+S_{\text{max}}))=1$.
@@
We provide additional remarks regarding the need for explicit bound specification of the log-time CDF and AULTC and the non-convexity of the average log discrepancy in Appendix A.

### v2

@@ 1 Introduction
The Third International Consensus Definitions for Sepsis (Sepsis-3) are the benchmark definitions for sepsis, the dysfunctional immune response to an infection, which are used in trial eligibility criteria [1] and phenotyping studies [2] . In the critical care literature, sepsis is seen as a heterogeneous disease, a pathway arrived upon from a multitude of infectious origins, anatomical sites, and comorbid profiles. There is increasing recognition of the importance of time in understanding sepsis and sepsis progression, as seen in early-warning systems [3] , variation in prediction utility [4] , and endotype characterization [5] .
Meanwhile, structured data streams, which are the data resources these tools use, are often incomplete records with insufficient information to render or confirm a diagnosis of sepsis [6]. A more complete alternative is the discharge summary, or more generally, the case report, in which a clinician documents the clinical findings pertinent to the case and care management. The impracticality of using the case report is that it is only written in full after the encounter is over, i.e., with a less granular timestamp than the events contained within. In principle, having a more complete, fine-grained temporal record could enable better temporal analyses of patients with sepsis, leading to better understanding and treatment improvements.
At present, the (semi-)public critical care repositories at the research frontier contain high resolution structured signals and low resolution textual signals. While several concerted efforts have focused on extracting temporal information, they have largely focused on temporal relations between clinical concepts [7] rather than on event timing. Those that have focused on timing [8, 9] used excerpted summaries and relatively small sample sizes.
In order to train sepsis models that can consider sources of information references in free-text and that have fine time-granularity, we seek to generate larger samples of textual time series. To do so, we introduce a novel Sepsis-3 textual time series corpus from the Pubmed Open Access (PMOA) Subset. We adopt the LLM-as-annotator approach, evaluating its ability to extract clinical findings and its ability to accurately associate them with timestamps. We augment existing clinical annotations with our own on PMOA, and contrast LLM annotations with our own and that of Frattallone et al. (2024) [9] . To provide checks on their performance at scale, we consider a larger subsample for inter-LLM annotator agreement, and show the performance levels across several model sizes, which provides valuable context for cost-based analyses. Finally, these assessments provide evidence of the LLM annotations quality and our pipeline, which are then applied to generate the textual time series corpus for sepsis (T2S2).
@@ 1.1 Related work
Numerous critical care and sepsis datasets are available (semi-)publicly [10, 11] , including several with de-identified clinical discharge summaries: MIMIC-III [12] and MIMIC-IV [13] . Sepsis phenotyping (including Sepsis-3) has been conducted using a variety of methods on structured data [14] , although structured data is often highly missing, leading to suboptimal phenotypes [15]. Additionally, the timing of sepsis phenotyping is crucial, in that delays in sepsis prediction can nullify predictiveness and lead to burdensome and unhelpful analytic models [4] .
Several works have used the i2b2 (a competition subset of MIMIC-II/III) to construct timelines of clinical concepts [8, 9] . In Leeuwenberg et al. [8] , the event spans are already given, and they focus on an excerpt of the discharge summary. Hence, we focus our direct comparison on that of Frattallone et al [9] . In contrast to these works, we adopt an annotation process focused on temporal assignment to clinical finding rather than i2b2 clinical concept [16] , which allows for greater specificity of the event in the absence of additional context. Additionally, as compared to Frattallone et al. [9] , we focus on using text alone for clinical time series construction, which is necessary for PMOA since no complementary data source is available.
Broadly in the medical domain, LLMs are showing promise in medical tasks such as summarization [17] . In fact, some findings suggest that medical fine-tuning rarely leads to improved model over their corresponding foundation models [18] . In our analysis, we primarily use foundation LLMs and show that their scale and instruction-tuning are central to their performance.
@@ 2 Methods
Figure 1: PMOA T2S2 pipeline (left) and Sepsis-3 confusion matrices (right)
Our task is the following: given a biomedical free-text corpus, construct a textual time series case report corpus of patients with Sepsis-3. We define a textual time series as a list of clinical findings, each with a timestamp (absolute or relative to case presentation) for an individual. A clinical finding is a free-text specification of an entity pertaining to or with the potential to affect the person’s health.
Compared to clinical concept annotations [16, 7] , we choose to expand the text span determination to increase the specificity of the meaning attached to the finding. We note the following distinctions of our notion of clinical finding with that of i2b2 clinical concept, as per the i2b2 guideline specifications [16] :
@@
The other design choice we make is to alter the time expression to disambiguate timing by capturing the start time of the perception of the clinical event. For example, if a patient has a “three-day history of fever”, the clinical event “fever” is assigned to -72 hours, since it is first perceived then. If the patient has a “history of smoking”, the start time is fully ambiguous, so the time of perception is the time that smoking is assessed, e.g., at the start of the encounter, and the event is “history of smoking”, rather than “smoking”. Our clinical finding is defined as a contiguous text span representing a temporally localizable patient event, adjusted for conjunction splitting and temporal perception.
@@ 2.1 Data
We used two data sources: (1) the absolute timeline annotations [9] , and (2) the Pubmed Open Access subset [19] (see Table 2.1.1).
@@
The absolute timeline annotations [9] came from physician annotations of 20 discharge summaries following the i2b2 concept and annotation guidelines (note, the original competition annotation focused on discharge summary excerpts rather the entire note), with the addition of interval timestamping of the identified concepts. For our analysis, we used the lower bound of the interval (specified as a mean time) as it denotes the time when the finding may first be observed. We used only the probabilistic annotations, since the bounded annotation lower bound could precede the actual occurrence of the finding. Fifteen of the summaries come from i2b2 and five are from MIMIC IV (annotated similarly)—hence we call this dataset “i2m4”.
[…]
Dataset summary. Bold cells indicate contributed labeling Corpus i2m4 Sepsis-10   Sepsis-100 N (unlabeled) 310; 331,794 2,139 2,139 N (labeled) 20 10 100 Source i2b2; MIMIC IV PMOA PMOA Note type discharge summary case report case report Labeler clinicians clinicians O1-preview Phenotyper clinicians clinicians QwQ-32B-Preview Methods O1-preview O1-preview DeepSeek R1 IQ1 DeepSeek R1 IQ1 DeepSeek R1 IQ1 L3.3 70B Instruct L3.3 70B Instruct L3.3 70B Instruct L3.1 70B Instruct L3.1 70B Instruct L3.1 70B Instruct Mixtral 8x7B Mixtral 8x7B Mixtral 8x7B  Instruct v0.1  Instruct v0.1  Instruct v0.1 L3.1 8B Instruct L3.1 8B Instruct L3.1 8B Instruct Ablations No role No role No role Base: No conjunction No conjunction No conjunction L3.3 70B Instruct  instruction  instruction  instruction No few-shot prompting No few-shot prompting No few-shot prompting Sensitivity analyses With upper bound With upper bound With upper bound With upper bound + With upper bound + With upper bound +  i2b2 event typing  i2b2 event typing  i2b2 event typing
@@ 2.1.2 PMOA (sepsis-10, sepsis-100) Dataset
The second dataset comes from the Pubmed Open Access Subset, an open-access, free-text corpus of roughly 1.5 million biomedical publications [19] . Because the majority of these publications are not case reports, we built an extraction pipeline to recover critical care case reports, i.e., a clinical research informatics critical care cohort.
We used the PMOA Subset (version as of April 25, 2024). We removed the header and footer of the free-text which contained the Title, Abstract, and References sections, and retained the ==== Body section. To identify potential case reports, we required that the free-text match the case-insensitive regular expressions case (report|presenta) and year-? ?old. We compared the query output to PubMed case report metadata and found our regex approach had much higher specificity in a sample of 20 random reports.
To identify potential sepsis case reports, we filtered the potential case reports using the case-insensitive regular expressions (sepsi||septic) and (critical||intensive) care. Then, we used the an LLM query to extract sepsis cases, as well as the number of cases present, and the age and gender. We used GPT-4o and Llama-3.1-8B-Instruct for the phenotyping query screen. Reports were omitted if the number of cases identified was not equal to 1. As a screen, we chose to include the case report if either GPT-4o or Llama-3.1-8B-Instruct identified the report as having Sepsis-3.
@@ 2.2 Textual Time Series Annotations
18 years old $|$ 0
male $|$ 0
admitted to the hospital $|$ 0
fever $|$ -72
rash $|$ -72
acne $|$ -672
minocycline $|$ -672
increased WBC count $|$ 0
eosinophilia|| 0
systemic involvement|| 0
diffuse erythematous or maculopapular eruption|| 0
pruritis $|$ 0
DRESS syndrome $|$ 0
fever persisted $|$ 0
rash persisted $|$ 0
discharged $|$ 24
@@
For the i2m4 dataset, we also compared against a BERT-based baseline (BERT 5-CV), our adaption of the Frattallone et al. model (2024) [9] . In their work, they transformed the temporal prediction task into multiclass classification and used a multilayer perceptron (MLP) head with a softmax and cross-entropy loss for training. To predict relative time (hours from case presentation), we made the following adaptations. Similarly to the classifier module, we used a 3-layer MLP regressor head from the output of the BERT event span features and output the prediction means and standard deviations. The regression task (predicting the lower bound) was optimized using L1 loss. We added the BERT features SUBJ_START and SUBJ_END tokens surrounding the event span [20] and removed the dropout layers inside the MLP heads, leaving only one at the final layer. Regarding the timestamps used for prediction, while many events occurred near the admission time, a subset extended months or years away, complicating the design of a regressor capable of both spanning wide magnitudes and maintaining precision for the majority of the means close to the admission time. To address this, we filtered the findings by their magnitudes, retaining them if the means were within two weeks of the admission time and the standard deviations were under two weeks. Filtering reduced the dataset size (in number of findings) from 3,472 to 3,350. Training was carried out for 50 epochs with batch size 32 and learning rate 3e-5. The model training used 5-fold cross-validation.
@@ 2.3 Assessments
We assessed the textual time series with the following methods: event match rate, concordance (c-index), median absolute time discrepancy, and log-time discrepancy. To determine event matches between two lists of strings, one for the label clinical findings and one for the predicted clinical findings, we used a recursive best match procedure, where a match is identified by the distance between two strings (see Appendix 2 for pseudocode). Specifically, the recursive best match was used to reduce erroneous surjective matches, i.e., multiple events from the predicted event list mapping to the same event in the reference list. For each reference event, if multiple predicted events matched, the lowest distance match was selected (with event ordering used to break distance ties), and both the predicted and referent events were removed from the lists for the recursive call to match events. Then all matched events from the recursive calls were returned. To compute distance, we tried (1) Levenshtein distance, (2) cosine distance between mean sentence embeddings of BERT (bert-base-uncased) and (3) cosine distance between mean sentence embeddings of PubMedBert (S-PubMedBert-MS-MARCO) [21]. Upon manual review of event matches, we selected option 3 with a cosine similarity cutoff of 0.1.
@@ 2.3.1 Log-Time Cumulative Distribution Function
We define the log-time cumulative distribution function as follows. For each matched event $i$, compute the discrepancy $s_{i}=|t^{p}_{i}-t^{r}_{i}|$. then truncate at the maximum cutoff $S_{\text{max}}$ (e.g., 1 year = 8760 hours): $s^{\prime}_{i}=\min(s_{i},S_{\text{max}})$.
Apply the log transformation: $x_{i}=\log(1+s^{\prime}_{i})$ Sort $x_{i}$ in ascending order: $x_{(1)}\leq x_{(2)}\leq\cdots\leq x_{(k)}$. The CDF $F(x)$ is defined as the fraction of log-transformed discrepancies less than or equal to $x$: $F(x)=\frac{1}{k}\sum_{i=1}^{k}\mathbf{1}_{\{x_{(i)}\leq x\}}$ where $\mathbf{1}$ is the indicator function. Then, we get the following boundary values: $F(0)\geq 0$ (depending on the smallest discrepancy) and $F(\log(1+S_{\text{max}}))=1$.
@@
Additional remarks on the necessity of explicit bounds for the log-time CDF and AULTC, as well as the non-convexity of the average log discrepancy, are provided in Appendix References.